# Efficient and Accurate Spatial Mixing of Machine Learned Interatomic Potentials for Materials Science

### v1

@@ 4.1 Constrained potential fitting
Figure 6: A schematic representation of the constrained fitting of a cheap potential to an expensive one. Top: A representation of the local potential energy surface of the expensive reference model (orange curve). Data has been sampled from this potential energy surface (crosses). Bottom: The cheap model, fit to this data in order to locally match the expensive model. It approximately matches the ‘soft’ constraint data (pink crosses) and exactly matches the ‘hard’ constraint data (green crosses).
4.1.1 Selecting constraints
The requirements of the cheap potential are split into two types: soft constraints (loose matching), and hard constraints (tight matching). Archetypal hard constraints are the elastic constants, to ensure seamless matching of the long range elastic stress fields between the cheap and expensive potentials in a solid state simulation [35].
@@ 4.1.1 Selecting constraints
$\min_{\mathbf{c},||\mathbf{A}_{\mathrm{H}}\mathbf{c}-\mathbf{y}_{\mathrm{H}}||^{2}<\alpha}\left(||\mathbf{y}_{\mathrm{S}}-\mathbf{A}_{\mathrm{S}}\mathbf{c}||^{2}\right)$ (1)
where regularization terms have been excluded for simplicity. The hard constraint is imposed by constraining the fit to the subspace $||\mathbf{A}_{\mathrm{H}}\mathbf{c}-\mathbf{y}_{\mathrm{H}}||^{2}<\alpha$, where $\alpha$ is the maximum allowed error for the model on the hard constraint data. Expressing this constraint through a Lagrange multiplier argument gives us the augmented Lagrangian
$\mathcal{L}(\mathbf{c},\lambda)=||\mathbf{y}_{\mathrm{S}}-\mathbf{A}_{\mathrm{S}}\mathbf{c}||^{2}+\lambda(||\mathbf{y}_{\mathrm{H}}-\mathbf{A}_{\mathrm{H}}\mathbf{c}||^{2}-\alpha).$ (2)
@@
$\min_{\mathbf{c}}(||\mathbf{y}_{\mathrm{S}}-\mathbf{A}_{\mathrm{S}}\mathbf{c}||^{2}+\lambda||\mathbf{y}_{\mathrm{H}}-\mathbf{A}_{\mathrm{H}}\mathbf{c}||^{2}).$ (3)
@@
From this expression, it is clear that the quadratic constraint is equivalent to adding the hard constraint configurations to the fit with weights scaled by $\lambda$. Obtaining a solution to the full constrained problem therefore requires finding the minimum $\lambda$ such that the solution to (3) lies on the surface of the constraint subspace $||\mathbf{A}_{\mathrm{H}}\mathbf{c}-\mathbf{y}_{\mathrm{H}}||^{2}=\alpha$.
@@
Scripts to perform constrained linear fitting for ACE and UF3 potentials are provided in the accompanying code [57]. To perform the $\lambda$ hyperparameter search, the linear problem is solved repeatedly. Initially, $\lambda$ is increased logarithmically until $||\mathbf{A}_{\mathrm{H}}\mathbf{c}-\mathbf{y}_{\mathrm{H}}||^{2}<\alpha$. Interval bisection is then used to find $\lambda$ such that $||\mathbf{A}_{\mathrm{H}}\mathbf{c}-\mathbf{y}_{\mathrm{H}}||^{2}-\alpha<\mathrm{tol}$, where $\mathrm{tol}$ is set by default as $\alpha/10$. $\mathbf{A}_{\mathrm{H}}$ and $\mathbf{A}_{\mathrm{S}}$ only need to be assembled once, and as the cheap potentials are small (often $\lesssim 100$ parameters), the whole constrained fitting process takes approximately 5 minutes on a single CPU core.
4.1.2 Cheap potential data
@@
In each of the three case-studies, regions far from the defect cores in each domain resembled near-bulk high temperature MD. Soft constraint data was generated to match this. For Si a 4096-atom diamond structure bulk cell was simulated at 500 K for 10 ps, with snapshots saved every 0.5 ps. For W and Fe, 2000-atom body centered cubic (BCC) bulk cells were simulated at 1200 K, with snapshots saved every 0.1 ps. The soft constraints consisted of the total energy and forces from each snapshot.
For the hard constraints, following [38], data was generated by applying multiple shear and expansion deformations to $5\times 5\times 5$ bulk supercell of Si, Fe and W lattices and saving the resulting energies. 17 configurations were generated between ±$\pm 0.5\%$ strain for all of the following strain states: hydrostatic, uniaxial [100], uniaxial [111], shear ([100], (010)), shear ([110], (001)) and shear ([110], ($\bar{1}$2)).
@@
For the Si and Fe ACE potentials, an $\alpha$ value was selected that corresponded to an enforced energy error of less than $10^{-7}$ eV/atom on the hard constraint data. For Si, the corresponding $\lambda$ value was $3.25\times 10^{9}$. For Fe, $\lambda$ was found to be $5.5\times 10^{9}$. For the W potential the constrained fitting process failed due to the UF3 potential not being capable of simultaneously fitting the hard constraint data and maintain a low RMSE on the soft constraints. By searching the space of $\lambda$ values, it was found that above approximately $\lambda$ = $4.3\times 10^{5}$, $||\mathbf{A}_{\mathrm{H}}\mathbf{c}-\mathbf{y}_{\mathrm{H}}||^{2}$ stopped decreasing. $\lambda$ was therefore set to $4.3\times 10^{5}$.
@@ 4.2 Region tracking
Figure 7: Region construction schematic. From left to right: Starting with a cell containing an initial defect, it is identified either using a user-specified LAMMPS group or by the output vector from a user-defined LAMMPS fix. The defect atoms are then iterated over, with atoms that sit less than one core radius out from any defect atom being added to the core region. This construction process is then repeated for the atoms in the blended region.
@@
In order to satisfy all three of these simultaneously, the region building and tracking were implemented directly in LAMMPS through a fix. By re-using the same neighbour lists that are built for the pair style evaluation step, this fix can quickly rebuild regions. Additionally, the fix is MPI local; whilst a limited amount of communication between neighbouring processes is necessary, no expensive global communication is required.
@@
4.2.1 Seed atom identification
Seed atoms are identified in one of two ways; through a LAMMPS group or through querying the output vector of a separately defined LAMMPS fix. Tracking through a predefined LAMMPS group is for situations where seed atoms are unchanging, and are known at the start of the simulation, as is the case with He atoms in W. Tracking through the output vector of a user-supplied fix is useful for tracking defects which are not tied to any one particular set of atoms, for example a vacancy, dislocation or the Fe dumbbell interstitial described in the previous section. Seed atom identification is schematically represented in the first two panels of Fig. 7.
4.2.2 Region construction
@@
During region construction two arrays are populated, i2_potential and d2_eval. Both arrays are of size number of potentials $\times$ number of atoms. The i2_potential array holds integers and is used to determine which atoms are to be evaluated with each potential; i2_potential[1][10] = 1 indicates that atom 10 should be evaluated with potential 1, whilst i2_potential[1][10] = 0 indicates that it should not be. d2_eval holds double-precision real numbers and describes how much each potential contributes to forces on atoms. By allowing d2_eval to hold fractional values, it is possible to introduce a blending region in which atoms move under a mixture of the forces from the two separate potentials; d2_eval[0][10] = 0.4 and d2_eval[1][10] = 0.6 indicates that atom 10 should get 40% of its’ forces from potential 0 and 60% from potential 1.
@@
6:     for j $\leftarrow$ neighbours of atom i do
7:         if $||\mathrm{pos(i)}-\mathrm{pos(j)}||$ $<$ r_core then
@@
2:     for j $\leftarrow$ neighbours of atom i do
3:         if $||\mathrm{pos(i)}-\mathrm{pos(j)}||$ $<$ r_blend then
@@
2:     for j $\leftarrow$ neighbours of atom i do
3:         if $||\mathrm{pos(i)}-\mathrm{pos(j)}||$ $<$ r_buffer then
@@ 4.3 Spatial potential mixing
The mixing scheme implemented in ML-MIX is force-mixing, which is a scheme commonly used in QM/MM. For in-depth reviews of force-mixing in the QM/MM context, please see Bernstein et al [35]. Here we briefly review the force-mixing method, and discuss simplifications that can be made when using two local potentials.
4.3.1 Force mixing
Figure 8: An illustration of force-mixing with interatomic potentials. A mixed domain is split into two non-interacting subdomains, which are composed of all the atoms in the core and blending regions for each potential (solid line), plus an extra buffer region containing atoms up to the potential cutoff radius out (dotted line). Partial force vectors $\mathbf{F}_{\mathrm{c}}$ and $\mathbf{F}_{\mathrm{e}}$ are obtained by evaluating the cheap and expensive subdomains respectively. These partial vectors are then combined into the full force vector on the system $\mathbf{F}$. In this process, forces on buffer atoms are discarded and forces on blending atoms are mixed between potentials.
@@ 4.3.1 Force mixing
$\begin{split}\Lambda_{k\text{-buffer}}=\bigcup\Big{\{}i\in\Lambda\setminus\Lambda_{k}\,\Big{|}\,\exists j\in\Lambda_{k},\\ \text{ such that }d(i,j)\leq r_{\mathrm{cutoff}}\Big{\}}.\end{split}$ (5)
Note that through the dependence of local energies on local atomic positions, there is an implicit dependence on atomic positions up to two cutoff radii out. Hence, r_buffer need not be any greater than the potential cutoff.
@@
5:              Remove j from neighbour list
@@
11:     Restore neighbour list
@@ 4.4 Si vacancy migration
To generate the vacancy structure, a 10$\times$10$\times$10 (8000) atom Si super-cell was generated, and the atom nearest the centre was removed. Initial guesses for the minimum energy pathway were formed by linearly interpolating the migration of a neighbouring atom into this vacancy. Including the start and end states, 9 images were generated along this pathway, which were then relaxed using the LAMMPS NEB package. For the ML/ML simulation, the seed atoms for the expensive potential region were a union of all atoms immediately surrounding the vacancy in the final and end states. These neighbours were identified using coordination analysis with a cutoff of 4.0 Å, with all atoms that have 15 neighbours being selected. An additional 4.0 Å of core region was added around these atoms, as well as 4.0 Å of blending region. 6.0 Å of buffer region was then included, which matches the cutoff of the cheap and expensive potentials. Region rebuilding was switched off, this is because dynamically changing the expensive potential region during a relaxation will often prevent convergence. For both the all-expensive and ML/ML NEB simulations, no energy tolerance was set (as total energies are not present in the ML/ML simulation) and the force tolerance was set such that $\|\mathbf{F}\|_{\infty}<0.0005$ eV/Å. Tighter tolerances than this would frequently fail to converge for the ML/ML simulation. If tighter tolerances are necessary, then we recommend switching to the all-expensive regime for the last steps of a relaxation.
@@ 4.5 Si stretched bond
To generate the structure, a $10\times 10\times 10$ (8000 atom) periodic supercell of bulk Si was constructed and a bond in the centre of the structure was stretched an additional 0.1 Å along its length and fixed using the rigid package in LAMMPS [61]. This block was then pre-thermalised to 300 K in a 2 ps simulation with a Langevin thermostat using only the expensive potential. This thermalised initial structure was used as the starting point for all subsequent comparison simulations, in which the average force on this bond was recorded over 100 ps. For the all-expensive simulations, an $NVE$ ensemble was used. For the ML/ML simulations, the seed atoms were selected to be the atoms involved in the stretched bond, no blending region was used (abrupt force mixing), as well as a weak thermostat of time constant 1.0 ps. For all simulations a buffer region size of 6.0 Å was used (which matches the cutoff of both potentials). The size of the core region was varied between simulations, as described in section. 2.3. The expensive potential region was rebuilt about this seed atom every timestep.

### v2

@@ 4.1 Constrained potential fitting
Figure 8: A schematic representation of the constrained fitting of a cheap potential to an expensive one. (a): A representation of the local potential energy surface of the expensive reference model (orange curve). Data has been sampled from this potential energy surface (crosses). (b): The cheap model, fit to this data in order to locally match the expensive model. It approximately matches the ‘soft’ constraint data (pink crosses) and exactly matches the ‘hard’ constraint data (green crosses).
Figure 9: Automated selection of lambda value during a single constrained fit to satisfy a given alpha. The vertical axis shows the energy RMSE per atom. The target alpha value is shown as a dashed red horizontal line. Each blue point represents a step on the search.
We first consider the challenge of selecting suitable constraints. The requirements of the cheap potential are split into two types: soft constraints (loose matching), and hard constraints (tight matching). Archetypal hard constraints are the elastic constants, to ensure seamless matching of the long range elastic stress fields between the cheap and expensive potentials in a solid state simulation [Bernstein_2009].
@@
[…]
where regularization terms have been excluded for simplicity. The hard constraint is imposed by constraining the fit to the subspace $\mathbf{A}_{\mathrm{H}}\mathbf{c}-\mathbf{y}_{\mathrm{H}}||^{2}<\alpha$, where $\alpha$ is the maximum allowed error for the model on the hard constraint data. Expressing this constraint through a Lagrange multiplier argument gives us the augmented Lagrangian
[…]
@@
[…]
@@
From this expression, it is clear that the quadratic constraint is equivalent to adding the hard constraint configurations to the fit with weights scaled by $\lambda$. Obtaining a solution to the full constrained problem therefore requires finding the minimum $\lambda$ such that the solution to (3) lies on the surface of the constraint subspace $\mathbf{A}_{\mathrm{H}}\mathbf{c}-\mathbf{y}_{\mathrm{H}}||^{2}=\alpha$.
@@
Scripts to perform constrained linear fitting for ACE and UF3 potentials are provided in the accompanying code [ML-MIX]. To perform the $\lambda$ hyperparameter search, the linear problem is solved repeatedly. Initially, $\lambda$ is increased logarithmically until $\mathbf{A}_{\mathrm{H}}\mathbf{c}-\mathbf{y}_{\mathrm{H}}||^{2}<\alpha$. Interval bisection is then used to find $\lambda$ such that $\mathbf{A}_{\mathrm{H}}\mathbf{c}-\mathbf{y}_{\mathrm{H}}||^{2}-\alpha<\mathrm{tol}$, where $\mathrm{tol}$ is set by default as $\alpha/10$. An example $\lambda$ search to satisfy a constraint of $\alpha=10^{-7}$ is shown in Fig. 9. $\mathbf{A}_{\mathrm{H}}$ and $\mathbf{A}_{\mathrm{S}}$ only need to be assembled once, and as the cheap potentials are small (often $\lesssim 100$ parameters), the whole constrained fitting process takes approximately 5 minutes on a single CPU core.
4.2 Generation of synthetic data
Figure 10: Demonstration of convergence of key quantities with hard constraint strength for the Si ACE 2, 10 cheap model. Error is shown compared to the Si ACE 4, 20 expensive model. In each case, the unconstrained case is shown by a dashed horizontal red line and the result at each constraint strength ($\alpha$) is shown by a blue point. (a): The L2 norm of the error over $\mathrm{C}_{11}$, $\mathrm{C}_{12}$ and $\mathrm{C}_{44}$. (b): The absolute error in lattice parameter. Note that at a constraint strength of $10^{-7}$, the lattice parameter error was 0 up to the number of decimal places measured (5).
@@ 4.2 Generation of synthetic data
Simple high-temperature bulk equilibrium dynamics was selected as the soft constraint data. For Si, a 4096-atom diamond structure bulk cell was simulated at 500 K for 10 ps, with snapshots saved every 0.5 ps. For W and Fe, 2000-atom body centered cubic (BCC) bulk cells were simulated at 1200 K, with snapshots saved every 0.1 ps. The soft constraints consisted of the total energy and forces from each snapshot.
[…]
@@
For the Si and Fe ACE potentials, an $\alpha$ value was selected that corresponded to an enforced energy error of less than $10^{-7}$ eV/atom on the hard constraint data. For Si, the corresponding $\lambda$ value was $3.25\times 10^{9}$. For Fe, $\lambda$ was found to be $5.5\times 10^{9}$. To a reader familiar with MLIPs, a target per-atom-energy error of $\alpha=10^{-7}$ eV/atom may seem very small (a standard aim when fitting MLIPs is a per-atom-energy error of around 1 meV). However, we found that pushing $\alpha$ to such small values was necessary to achieve the tight property matching that we were aiming for. This is demonstrated for the Si cheap potential in Fig. 10; such small $\alpha$ values were necessary to match elastic constants within a fitting error of $\sim 2$ GPa and lattice parameter within a fitting error of $10^{-5}$ Å. For the W potential the constrained fitting process failed due to the UF3 potential not being capable of simultaneously fitting the hard constraint data and maintain a low RMSE on the soft constraints. By searching the space of $\lambda$ values, it was found that above approximately $\lambda$ = $4.3\times 10^{5}$, $\mathbf{A}_{\mathrm{H}}\mathbf{c}-\mathbf{y}_{\mathrm{H}}||^{2}$ stopped decreasing. $\lambda$ was therefore set to $4.3\times 10^{5}$.
4.3 Additional DFT data
In the case of the cheap and expensive potential fit for the W dislocation and He – W implantation studies, additional DFT data was acquired. The DFT parameters used matched those described in [Nutter_2024_supp]. For the dislocation mobility simulations, the dataset in [Nutter_2024_supp] was extended to include 1000 K surface and dislocation kink configurations. 40 structures for both (110) and (112) surfaces were generated, each containing 45 and 48 atoms, respectively. 12 structures for each of the two kink types (left and right) were generated, containing 330 and 345 atoms, respectively. For the He – W implantation study, 129 snapshots of small 91-atom implantation cells were generated, from simulations of He implantations at 10 eV and 100 eV. These were added to the full He – W dataset provided by Nutter et al [Nutter_2024], which was then used for generation of the expensive and cheap potentials for the He – W implantation simulation.
@@ 4.4 Region tracking
Figure 11: Region construction schematic. From left to right: Starting with a cell containing an initial defect, it is identified (orange atoms in panel two) either using a user-specified LAMMPS group or by the output vector from a user-defined LAMMPS fix. The defect atoms are then iterated over, with atoms that sit less than one core radius out from any defect atom being added to the core region. This construction process is then repeated for the atoms in the blended region.
@@
In order to satisfy all three of these simultaneously, the region building and tracking were implemented directly in LAMMPS through a custom fix. By re-using the same neighbor lists that are built for the pair style evaluation step, this fix can quickly rebuild regions. Additionally, the fix is MPI local; whilst a limited amount of communication between neighboring processes is necessary, no expensive global communication is required.
@@
Seed atoms are identified in one of two ways; through a LAMMPS group or through querying the output vector of a separately defined LAMMPS fix. Tracking through a predefined LAMMPS group is for situations where seed atoms are unchanging, and are known at the start of the simulation, as is the case with He atoms in W. Tracking through the output vector of a user-supplied fix is useful for tracking defects which are not tied to any one particular set of atoms, for example a vacancy, dislocation or an Fe dumbbell interstitial. Seed atom identification is schematically represented in the first two panels of Fig. 11.
@@
During region construction two arrays are populated, i2_potential and d2_eval. Both arrays are of size number of potentials $\times$ number of atoms. The i2_potential array holds integers and is used to determine which atoms are to be evaluated with each potential; i2_potential[1][10] = 1 indicates that atom 10 should be evaluated with potential 1, whilst i2_potential[1][10] = 0 indicates that it should not be. d2_eval holds double-precision real numbers and describes how much each potential contributes to forces on atoms. By allowing d2_eval to hold fractional values, it is possible to introduce a blending region in which atoms move under a mixture of the forces from the two separate potentials; d2_eval[0][10] = 0.4 and d2_eval[1][10] = 0.6 indicates that atom 10 should get 40% of its’ forces from potential 0 and 60% from potential 1. Two possible blending functions are implemented in ML-MIX - linear blending and cubic blending. A brief discussion of the impact on energy conservation of these two blending schemes is presented in section S4 of the supplementary information.
@@
6:  for j $\leftarrow$ neighbors of atom i do
7:   if $\mathrm{pos(i)}-\mathrm{pos(j)}||$ << r_core then
@@
2:  for j $\leftarrow$ neighbors of atom i do
3:   if $\mathrm{pos(i)}-\mathrm{pos(j)}||$ << r_blend then
@@
2:  for j $\leftarrow$ neighbors of atom i do
3:   if $\mathrm{pos(i)}-\mathrm{pos(j)}||$ << r_buffer then
@@ 4.5 Spatial potential mixing
The mixing scheme implemented in ML-MIX is force-mixing, which is a scheme commonly used in QM/MM. For in-depth reviews of force-mixing in the QM/MM context, please see Bernstein et al [Bernstein_2009]. Here we briefly review the force-mixing method, and discuss simplifications that can be made when using two local potentials.
Figure 12: An illustration of force-mixing with interatomic potentials. A mixed domain is split into two non-interacting subdomains (blue and orange), which are composed of all the atoms in the core and blending regions for each potential (solid line), plus an extra buffer region containing atoms up to the potential cutoff radius out (dotted line). Partial force vectors $\mathbf{F}_{\mathrm{c}}$ and $\mathbf{F}_{\mathrm{e}}$ are obtained by evaluating the cheap and expensive subdomains respectively. These partial vectors are then combined into the full force vector on the system $\mathbf{F}$. In this process, forces on buffer atoms are discarded and forces on blending atoms are mixed between potentials.
@@
$\begin{split}\Lambda_{k\text{-buffer}}=\Big\{i\in\Lambda\setminus\Lambda_{k}\,\Big|\,\exists j\in\Lambda_{k}\\ \text{ and }d(i,j)\leq r_{\mathrm{cutoff}}\Big\}.\end{split}$ (5)
Note that through the dependence of local energies on local atomic positions, there is an implicit dependence on atomic positions up to two cutoff radii out. Hence, r_buffer need not be any greater than the potential receptive field, as at this radius the forces match what they would be in non-mixed simulations on each domain. For MLIP architectures with a large receptive field due to message passing (e.g, MACE), one often does not need a buffer that captures it in its entirety; force errors reduce in size rapidly with buffer radius.
@@
5:     Remove j from neighbor list
@@
11:  Restore neighbor list
@@ 4.6 Si vacancy migration
To generate the vacancy structure, a $\times10$$\times$10 (8000) atom Si supercell was generated, and the atom nearest the centre was removed. Initial guesses for the minimum energy pathway were formed by linearly interpolating the migration of a neighboring atom into this vacancy. Including the start and end states, 9 images were generated along this pathway, which were then relaxed using the nudged-elastic-band (NEB) method [Henkelman_2000] as implemented in LAMMPS. Further analysis was conducted in Python using ASE [ASE]. For the ML/ML simulation, the seed atoms for the expensive potential region were a union of all atoms immediately surrounding the vacancy in the final and end states. These neighbors were identified using coordination analysis with a cutoff of 4.0 Å, with all atoms that have 15 neighbors being selected. An additional 4.0 Å of core region was added around these atoms, as well as 4.0 Å of blending region. 6.0 Å of buffer region was then included, which matches the cutoff of the cheap and expensive potentials. Region rebuilding was switched off; provided expensive atoms do not move far this allows relaxations to progress smoother and faster by stopping atoms changing the proportion of forces they gain from each potential as the minimization progresses. For both the all-expensive and ML/ML NEB simulations, no energy tolerance was set (as total energies are not present in the ML/ML simulation) and the force tolerance was set such that $\|\mathbf{F}\|_{\infty}<0.0005$ eV/Å. Tighter tolerances than this would frequently fail to converge for the ML/ML simulation. If tighter tolerances are necessary, then we recommend switching to the all-expensive regime for the last steps of a relaxation. The final energies were obtained through one-shot energy evaluations of the relaxed structures with the expensive potential.
@@ 4.7 Si stretched bond
To generate the structure, a $10\times 10\times 10$ (8000 atom) periodic supercell of bulk Si was constructed and a single bond was stretched an additional 0.1 Å along its length and fixed using the rigid package in LAMMPS [Kamberaj_2005]. This block was then pre-thermalized to 300 K in a 2 ps simulation with a Langevin thermostat using only the expensive potential. This thermalized initial structure was used as the starting point for all subsequent comparison simulations, in which the average force on this bond was recorded over 100 ps. For the all-expensive simulations, an $NVE$ ensemble was used. For the ML/ML simulations, the seed atoms were selected to be the atoms involved in the stretched bond, no blending region was used (abrupt force mixing), as well as a weak thermostat of time constant 2.0 ps applied to only the cheap potential region in order to stop temperature drift due to the ML/ML boundary . For all simulations a buffer region size of 6.0 Å was used (which matches the cutoff of both potentials). The size of the core region was varied between simulations, as described in section. 2.3. The expensive potential region was rebuilt about this seed atom every timestep. To compute the average bond force, samples of atomic forces were taken every 15 fs, long enough to avoid significant correlation between samples.